A firm expects to sell 26,000 units of its product at $11 per unit. Pretax income is predicted to be $61,000. If the variable costs per unit are $6, total fixed costs must be:

Multiple Choice

  • $69,000.

  • $156,000.

  • $225,000.

  • $43,000.

  • $286,000
Answer #1

Contribution margin=Sales-Variable costs

=(11-6)=$5 per unit

Total Contribution margin=(5*26,000)=$130,000

Contribution margin-Fixed costs=Target pre-tax income

130,000-Fixed costs=61,000

Fixed costs=130,000-61,000

=$69,000.